OpenStudy (anonymous):

i will give metal and fan.... Solve. 3 = -(-y + 6) A. -9 B. -3 C. 3 D. 9

OpenStudy (anonymous):

3 = –(–y + 6) 3 = y - 6 9= y

OpenStudy (mrbaconqwerty):

I think b, but im not sure

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

the question marks are negative sighns

OpenStudy (anonymous):

yeah idk why its like that

OpenStudy (anonymous):

\[3 = –(–y + 6)\] \[3 = y - 6\] \[9=6\]

OpenStudy (anonymous):

one negative sign for all the question marks idk eather

OpenStudy (anonymous):

sorry 9 = x not 6

OpenStudy (anonymous):

do you get it tho

OpenStudy (anonymous):

thanks yeah i get it